You add 3.00 g of Pb(NO3)2 (M.W. = 331.2 g/mole) to 75.0 mL of 0.100 M HCl, leading to formation of 0.873 g PbCl2 (M.W. = 278.1 g/mole) as a precipitate.


a) Which molecule is the limiting reactant?


 

b) What is the percent yield of PbCl2



1
Expert's answer
2021-12-20T09:12:03-0500

The balanced equation is below:

Pb(NO3)2 + 2HCl --> PbCl2 + 2HNO3

a)

"n(Pb(NO_3)_2)=3.00\ g\times\frac{1\ mol}{331.2\ g}=0.009058\ mol"


"n(HCl)=0.0750\ L\times\frac{0.100\ mol}{1\ L}=0.00750\ mol"

According to the equation, 1 mole of Pb(NO3)2 reacts with 2 moles of HCl. Actually, there are fewer moles HCl than Pb(NO3)2, hence HCl is the limiting reactant.


b)

"Theoretical\ yield\ PbCl_2=0.00750\ mol(HCl)\times\frac{1\ mol(PbCl_2)}{2\ mol(HCl)}\times\frac{278.1\ g(PbCl_2)}{1\ mol(PbCl_2)}=1.043\ g"


"\%\ yield=\frac{actual\ yield}{theoretical\ yield}\times100\%=83.7\%"


Answer: 83.7%